OpenStudy (anonymous):

Let f(x)=b^x be an exponential function such that f '(0)=1 Find b

OpenStudy (anonymous):

by f' you mean the derivative yes?

OpenStudy (anonymous):

yes, thanks

OpenStudy (anonymous):

Well, what is the derivative of b^x ?

OpenStudy (anonymous):

The value of b is e.

OpenStudy (anonymous):

how did you get there?

OpenStudy (anonymous):

It is a mathematical constant equal to 2.718(approx.) You must be knowing, I guess.

OpenStudy (anonymous):

I know what e is, how did you solve

OpenStudy (anonymous):

Did you take the derivative of f ?

OpenStudy (zarkon):

\[f'(x)=b^x\ln(b)\]

OpenStudy (anonymous):

Take log on both sides and then differentiate.

OpenStudy (anonymous):

You'll get the equation posted by Zarkon. Solve it by subs. x=0.

OpenStudy (anonymous):

If you had found the derivative (as Zarkon has given) you could plug in for \[1 = b^0(ln\ b)\] And find that ln(b) = 1 Therefore b = e.

OpenStudy (anonymous):

Did you get it? Or do you want a detailed working of the problem?

OpenStudy (anonymous):

got it thanks

OpenStudy (anonymous):

Next time, don't be afraid to differentiate =)

OpenStudy (anonymous):

It's often not as bad as it seems.